The OpAmp configured below , the voltage gain, A, is given as:

WW

A. A = -Rf/R1
B. A = +Rf/R1
C. A =- (R1 +Rf/R1
D. A = 1 + Rf/R1
Answer #1

This is nothing but a inverting amplifier and we know voltage gain for this as -

A = V0/Vi = -Rf/R1